OpenStudy (anonymous):

the integral of (x/(4x^2+9)^2)

zepdrix (zepdrix):

Let's write it like this, it might help us to identify what's going on. \[\large \int\limits \frac{1}{(\color{orangered}{4x^2+9})^2}(\color{royalblue}{x\;dx})\] Let \(\large \color{orangered}{u=4x^2+9}\) Taking the derivative of our substitution gives us, \(\large du=8x\;dx\) Understand what's going on so far?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So if we take our du equation and divide each side by 8, we get,\[\large \color{royalblue}{\frac{1}{8}du=x\;dx}\] Do you see how these pieces will substitute in? I tried to color-code them to make it easier to see.

OpenStudy (anonymous):

yes makes sense

zepdrix (zepdrix):

\[\large \int\limits\limits \frac{1}{(\color{orangered}{4x^2+9})^2}(\color{royalblue}{x\;dx}) \qquad = \qquad \int\limits\limits \frac{1}{(\color{orangered}{u})^2}\left(\color{royalblue}{\frac{1}{8}du}\right)\] Do you understand how to solve it from here?

OpenStudy (anonymous):

\[1/8\int\limits_{?}^{?}(1/(u^2)\] what do i do after this

zepdrix (zepdrix):

We'll want to recall a rule of exponents: \(\large x^{-n}=\dfrac{1}{x^n}\) This works in reverse which is how we'll want to use it. \[\large \frac{1}{8}\int\limits \frac{1}{u^2}\;du \qquad=\qquad \frac{1}{8}\int\limits u^{-2}\;du\] And from here we want to apply the `Power Rule for Integration`. Remember how to do that?

OpenStudy (anonymous):

yes thank you so much for your help

OpenStudy (anonymous):

what answer did you come out with?

zepdrix (zepdrix):

Ummm let's see.\[\large = \frac{1}{8}\left(\frac{u^{-2+1}}{-2+1}\right)+C\] Which gives usssssss,\[\large -\frac{1}{8}u^{-1}+C \qquad = \qquad -\frac{1}{8u}+C\]From here you need to remember to undo your substitution. We want our final answer in terms of x, not u.

OpenStudy (anonymous):

thank you again. this really helped.

zepdrix (zepdrix):

cool, no problem c: